# Multi-User Chat Assistant (MUCA): a Framework Using LLMs to Facilitate Group Conversations

### 10 Jan 2024 v1

@@ 4.3.3 Quantitative Study with Small-size Groups
The quantitative comparisons for two chatbots are shown in Table 1. User engagement is compared with two metrics, i.e., the total words exchanged in the whole conversation (words ex.) and average message word length (Avg. msg len.). Participants tend to type and input more on average if a chatbot inspires them and engages them in a chat. Evenness evaluates the total letters a person types in an experiment, which is evaluated with the standard deviation (STD) of the length of messages. The consensus is obtained from the rates given by Group-A and Group-B, where the rate is represented by the number of agreements reached over the total number of tasks. A higher consensus rate means the chatbot can help the users to reach a consensus more efficiently.
@@
From Table 1, we can find that MUCA-advanced-small can help participants get better engagement in the conversation compared to MUCA-basic, although the evenness slightly decreases shown by a larger STD. It is also obvious that MUCA-advanced-small has a much higher consensus rate compared with MUCA-basic. From our observation and analysis, this is because: (1) MUCA-advanced-small gives more space for participants to discuss through fewer chime-in than MUCA-basic; (2) MUCA-advanced-small provides feasible suggestions for each party to directly help to reach agreement; Besides, it also provides insightful, un-repeated comments, and summary, etc. to indirectly help participants for efficient discussion. (3) MUCA-basic makes mistakes to re-transit the topic back to previous well-discussed topics and provides redundant information, which results in inefficient discussion.
4.3.4 Quantitative Study with Small-size and Medium-size Groups
Managing conversations in a medium-sized group typically presents more challenges compared to a small group setting. From the user perspective, as encouraging members to contribute evenly, the facilitator chatbot would be more effective in a medium-sized group. This is particularly important given that members in larger groups are prone to social loafing and free-riding behaviors. On the chatbot’s side, the increased participation of members in discussions increases the cognitive load required to organize and synthesize diverse opinions, making the management of larger groups significantly more difficult.
@@ 4.3.4 Quantitative Study with Small-size and Medium-size Groups
Additionally, we conducted a user study experiment involving 8 participants, focusing on Topic-A, and recorded its statistics in Table 1. We can observe that, despite an increase in group size, the performance of MUCA-advanced-medium remains relatively consistent. It shows a similar number of total messages exchanged and an even greater average message length, demonstrating MUCA-advanced-medium’s effectiveness. The user attitudes towards MUCA-advanced-medium are also comparable to those observed with MUCA-advanced-small. This suggests that the proposed MUCA is capable of managing interactions in larger groups effectively.
@@ 4.3.5 Practical Values for Multi-user Chatbots
As can be found in Figure 10, participants evaluated the practical values of MUCA. It reveals a strong preference for users to use MUCA-advanced-small for decision-making scenarios. Additionally, a majority of users favor the use of a chatbot to enhance chatting efficiency and facilitate brainstorming sessions. Compared to MUCA-basic, MUCA-advanced-small and MUCA-advanced-medium are obviously more preferred by users for their group conversations. However, there are still a few situations in which people do not want the engagement of a multi-user chatbot. The reason might be that some people prefer to work independently when seeking solutions to problems, or they believe that the chatbot should be further improved to become a more mature tool.
@@ 5 Limitation and Future Work
5.1 MUS: Limitations and Ad-hoc Solutions
Similar to previous research, we also discovered that resembling human behavior is challenging for the simulator. Additionally, constructing a high-quality and specialized user simulator for a specific task can be a labor-intensive process Walker et al. (1997); Liu and Lane (2017). Implementing a user simulator presents several challenges: (1) Generating natural language utterances with an LLM-based user simulator is particularly difficult when the number of words is small. For instance, the minimum length of utterance ($l_{min}=1$) and maximum length of utterance ($l_{max}=10$) extracted from chat history are quite small. To address this, we boosted $l_{min}$, $l_{avg}$, and $l_{max}$ for each virtual attendee correspondingly and also adjusted the number of words for the role of questioner. (2) GPT-4 cannot consistently follow instructions to generate a valid virtual attendee ID for the subsequent turn to speak. Instead, it tends to predict the LLM agent to speak next, particularly when someone directly addresses the LLM agent in the previous turn. To mitigate this issue, we opt to randomly select the virtual attendee and their respective speaking role. (3) Virtual attendees suffer from maintaining the same dialog act (e.g. asking questions, pinging the LLM agent) for consecutive turns. This issue might be due to the nature of the generative model which focuses on predicting the next token. To address this issue, we introduce a cool-down mechanism for some dialog acts such as asking questions, pinging the LLM agent, and topic shifting.
@@ 5.2 Future Work
Component Orchestration: The proposed MUCA framework integrates several components, enabling actions such as "participation encouragement" and "initiative summarization". These components have been carefully designed, tuned, and ranked to provide a harmonious experience to the chat participants. It can be beneficial to explore an easy plug-and-play method for users to design and incorporate new components into the framework without intensive tuning. Such a feature could be important, as different conversation scenarios may require different chatbot personas.
Human-in-the-loop Feedback Iteration: Full user studies for feedback are costly and time-consuming. To continuously improve chatbot post-launch, it is useful to collect implicit and explicit user behavior signals. This data should be easily transformable for automatic or semi-automatic chatbot enhancements.
Rapidly Advancing AI Technologies: The proposed MUCA framework is based on the recent state-of-the-art LLMs, each with its unique style and best practices for prompting. It would be beneficial to investigate methods for updating the underlying AI models without the need for complete re-prompting or component orchestration.
@@
Multi-Chatbot Design: The study concentrates on multi-user and single-chatbot interactions, but examining scenarios with multiple users and chatbots, each having diverse personas, interacting could be intriguing. For instance, in cross-disciplinary meetings, chatbots could serve as hosts, minute-takers, or subject matter experts, offering insights to human participants as needed.
@@ A.1 MUCA Algorithm Full Implementation
The full algorithm is shown in Algorithm A.1, which allows immediate chime-in for MUCA to fulfill participant’s request by checking the key word @mubot matching in the last utterance.
@@
Figure 12: Data flow for Dialog Analyzer, which includes participant feature extractor based on statistic computation and full prompting for three LLM-based modules – sub-topic status update, utterance feature extractor, and accumulative summary update. The placeholders (shown as purple) in the prompting are filled by sub-topics from a Sub-topic Generator, conversation signals such as attendee names and utterances in the short-term context window. The generated outputs (sub-topic status, accumulative summary, and sub-topic being discussed, shown in yellow) will be fed back to the sub-topics status update and accumulative summary update for the next-round generation.
@@ A.2.1 Sub-topic Generator
To facilitate Sub-topic Generation, users are required to provide a topic and may optionally include an agenda, hints, and attendee roles. Figure 11 shows the data flow for the Generator. Full prompting of input prompt templates where the purple text is represented as placeholders and is replaced by user-input Topic, Hints, and Attendee Role to generate 3 sub-topics (Agenda is optional and not shown in the example). Sub-topic Generator executes once before the chat starts, and the sub-topics remain the same throughout the chat.
Figure 13: Data flow for direct chatting dialog act, which includes the user-input message, its full prompting, and MUCA’s responses. The placeholders (shown in purple) in the prompting are filled by user-input topic, previous accumulative summary, sub-topics status, short-term context window, sub-topics being discussed, and the last utterance. The generated responses are shown to different types of user-input messages corresponding to different rules in the prompting.
@@ A.2.2 Dialog Analyzer
Figure 12 shows the data flow for the Dialog Analyzer. Only the participants feature extractor sub-module is based on statistic computation and the rest of the three sub-modules (sub-topic status update, utterance feature extractor, and accumulative summary update) are based on LLM inference results. Complete input prompt templates for the three LLM-based sub-modules where the purple and yellow text are represented as placeholders are shown. The purple ones are replaced by sub-topics from the sub-topic generator, conversation signals such as attendee names and utterances in the current context window, and the yellow ones are replaced by the generated outputs (sub-topic status, summary, and current sub-topic) from other modules. The outputs of the Dialog Analyzer will be fed into the downstream Utterance Strategies Arbitrator module to select the suitable dialog act for the response generation.
@@
Sub-topic Status Update: Three statuses are defined (i.e. not discussed, being discussed and well discussed), which defines differently for different types of topics to avoid ambiguity. For example, for goal-oriented communication tasks (i.e., decision making, problem-solving, and open discussion), well-discussed refers to a situation where a consensus has been reached by the majority of participants, and the remaining participants do not have any conflicting suggestions. However, in the context of chit-chat communication, well-discussed implies that participants engage in conversations closely related and relevant to the sub-topic at hand.
@@
We have discovered that differentiating between being discussed and well discussed proves to be challenging, despite having precise definitions for each term. Therefore, to improve the precision and accurately update the status, the prompting utilizes the CoT style (as shown in Figure 12). This method initially returns the sub-topic IDs, followed by summarizing each participant’s opinion using generated IDs based on the current context window and previous summary. Subsequently, it determines whether the discussion meets the definition of well discussed according to the summary. Finally, the status is updated based on the results.
@@
Figure 14: Data flow for sub-topic transition dialog act, which includes hints for different trigger conditions, its full prompting, and MUCA’s responses. The placeholders (shown as purple) in the prompting are filled by the current sub-topic, next sub-topic candidates, previous summary, and short-term context window with a size of $N_{sw}$. The generated responses are shown, corresponding to different trigger conditions.
@@
Accumulative Summary Update: This sub-module updates the summary for each participant under different sub-topics of the past $N_{lw}$ utterances based on the current context window with a size of $N_{sw}$, current sub-topics being discussed from utterance feature extractor and previous summary, enabling build a memory into the MUCA system. The succinct summary is updated and stored with respect to each participant under each sub-topic, which is inspired by the idea (LangChain) that extracts entities from stored messages and only returns information about entities referenced in the current run. Instead of relevant message retrieval based on the user-input query to generate a response, we provide a succinct summary for downstream modules to save computation costs for query understanding and retrieval.
@@ A.2.3 Utterance Strategies Arbitrator
Direct Chatting: In Figure 13, we show that the MUCA responds adaptively to different types of requests: greeting/acknowledgment, out-of-scope request, statement/comments, and regular request. To avoid hallucination, we disallow the MUCA to make invalid assumptions beyond the user-input information, but reasonable reasoning based on chatting history is allowed (See the second example in the figure).
@@
Figure 16: Data flow for in-context chime-in dialog act, which includes user-input context, its full prompting, and MUCA’s responses. The placeholders (shown as purple) in the prompting are filled by user-input message, previous accumulative summary, previous sub-topics status, current sub-topics being discussed, and short-term context window with a size of $N_{sw}$. The generated responses are shown, corresponding to different user-input contexts.
Figure 17: Prompting for multi-user simulator (MUS). All placeholders (purple) are filled by our pre-defined settings or outputs from upstream modules and the LLM agent when generating the input prompts for LLM. In the prompt, we use speaking personalities to replace utterance traits for better interpretation of LLM.
Participation Encouragement: The proposed MUCA provides customized message based on the lurker’s chatting history and chime-in status, which is recorded and transferred to the hints for LLM based on the frequency of each participant’s chime-in for various sub-topics. We can see the responses are customized by different participants’ chatting statuses in Figure 15.
Figure 18: Two examples of group conversations between the multi-user simulator (MUS) and the proposed chatbot (MUCA).
@@
Sub-topics Transition: Figure 14 shows the MUCA transits sub-topic based on different trigger conditions. It tries to boost participants’ interest by providing insights for both scenarios, but it only asks to switch the sub-topic in the second scenario where most of the participants have discussed the current sub-topic. Introducing a new topic can improve participants’ engagement, particularly those who have remained silent (lurkers).

### 16 Feb 2024 v3

@@ 4.3.3 Quantitative Study in Small-size Groups
The quantitative comparisons for two chatbots are shown in Table 1. User engagement is compared with two metrics, i.e., the average number of words exchanged per conversation (Engt.-Words/Conv.) and the average number of words per utterance (Engt.-Words/Utt.). Evenness is assessed by calculating the sample standard deviation (STD) of the word count input by each participant, expressed as a percentage of the mean. The consensus is obtained from the rates given by two groups for small group or one group for medium group, where the rate is represented by the number of agreements reached over the total number of tasks. A higher consensus rate means the chatbot can help the users to reach a consensus more efficiently.
From the comparison in Table 1, MUCA-small helps participants get better engagement, shown by increased Engt.-Words/Conv. and Engt.-Words/Utt., which indicates that participants were more inclined to engage in extensive conversations and to compose longer utterances. MUCA-small enhances evenness in Topic A discussions with a lower STD while keeping similar evenness in Topic B with a comparable STD over Baseline-small.
@@
It is also obvious that MUCA-small has a much higher consensus rate compared with Baseline-small. From our observation and analysis, this is because: (1) MUCA-small gives more space for participants to discuss through fewer chime-in than Baseline-small; (2) MUCA-small provides feasible suggestions for each party to directly help to reach agreement; Besides, it also provides insightful, un-repeated comments, and summary, etc. to indirectly help participants for efficient discussion. (3) Baseline-small makes mistakes to re-transit the topic back to previous well-discussed topics and provides redundant information, which results in inefficient discussion.
Additionally, Table 1 shows average scores from Group-A and Group-B on Efficiency, Conciseness, and Usefulness. For Topic-A, MUCA-small outperforms Baseline-small with 12.5%, 40.6%, and 28.1% higher scores, respectively. In Topic-B, MUCA-small scores slightly higher. The Overall Rating provided by participants also reflects similar trends: MUCA-small surpasses Baseline-small by 31.9% in Topic-A and 11.1% in Topic-B.
4.3.4 Quantitative Study in Small-size and Medium-size Groups
Managing conversations in medium-sized groups is more challenging than in small groups. A facilitator chatbot is supposed to be more effective in medium-sized groups, as it promotes even contribution among participants, countering social loafing and free-riding behaviors, which are common in larger groups. However, this increased participation raises the chatbot’s cognitive load for organizing and synthesizing diverse opinions, making larger group management more complex.
@@ 4.3.4 Quantitative Study in Small-size and Medium-size Groups
Additionally, we conducted a user study experiment in a medium group using Topic-A, and recorded its statistics in Table 1. We can observe that, MUCA-medium remains consistent performance despite the increased group size. It shows a comparable average number of words per utterance, and the user evaluation scores are comparable to those observed in MUCA-small, indicating MUCA-medium’s capability in effectively managing larger group interactions.
@@ 4.3.5 Practical Values for Multi-user Chatbots
As can be found in Figure 10, participants evaluated the practical values of MUCA. It reveals a strong preference for users to use MUCA-small for decision-making scenarios. Additionally, most users also favor the chatbot’s assistance in improving chat efficiency and aiding in brainstorming. Compared to Baseline-small, MUCA-small and MUCA-medium are more preferred by users for their group conversations. However, some participants still opt out of chatbot involvement. This reluctance may stem from a preference for independent problem-solving or a belief that chatbots need further development to be fully effective in serving as a mature tool.
@@ 5 Limitation and Future Work
There has been recent a surge of interest in LLMs due to their unprecedented capabilities that can be applied to several fields. However, LLMs do see many challenges, such as bias, fairness, toxicity, etc., and we refer readers to the numerous studies that are dedicated to addressing these problems. We will use this section to discuss issues that are particularly relevant to MUCA and MUS.
Multi-user Chat Assistant (MUCA): The proposed MUCA is a pioneering work dedicated to multi-user chats. Although MUCA is by no means a comprehensive solution, it provides significant insights that could pave the way for future work in this field. Our investigation has identified several challenges that call for further research. Here are a few of these challenges:
  • Firstly, MUCA encompasses seven sub-modules dedicated to dialog acts, but only the top-ranked one is chosen at a time for generating a response. This approach overlooks the potential to validate the response’s quality, as it is delivered irrespective of its merit. We believe that by requesting all the dialog act sub-modules to generate a response concurrently, MUCA will be able to comprehensively evaluate and validate all the response candidates. The final augmented response could then be synthesized by either selecting or merging from this pool of response candidates through another post-dialog-act procedure.
  • Secondly, in our user study cases, we adjusted the hyper-parameters ($N_{exe}$, $N_{sw}$, $N_{lw}$, $W$, $C$, $f$ and $g$) in MUCA based on experimental results on small to medium groups. For larger conversation groups, the effectiveness of the selected hyper-parameters needs empirically validation. Also, an automated mechanism determining these parameters based on the configurations and the environmental variables of the conversations can also greatly alleviate the burden of tuning these parameters.
  • Thirdly, compute resources requested by LLMs inference pose a significant constraint for MUCA, especially for large chat groups. To mitigate this challenge, we have slightly increased the execution interval ($N_{exe}$), which occasionally results in MUCA missing optimal opportunities for user interaction at the most suitable moment. Moreover, we have sometimes observed an interesting phenomenon wherein multiple participants simultaneously express the desire to directly engage with MUCA, leading to a surge in computational demands. How to handle high volume of LLM calls with limited compute resources, while simultaneously striving to preserve the responsiveness of MUCA to the best extend, is a topic that worth further investigation.
Multi-user Simulator (MUS): Similar to previous research, we also discovered that resembling human behavior is challenging for the simulator. Additionally, constructing a high-quality and specialized user simulator for a specific task can be a labor-intensive process Walker et al. (1997); Liu and Lane (2017). Implementing a user simulator presents several challenges:
  • Firstly, generating natural language utterances with an LLM-based user simulator is particularly difficult when the number of words is small. For instance, the minimum length of utterance ($l_{min}=1$) and maximum length of utterance ($l_{max}=10$) extracted from chat history are quite small. To address this, we boosted $l_{min}$, $l_{avg}$, and $l_{max}$ for each virtual attendee correspondingly and also adjusted the number of words for the role of questioner.
  • Secondly, LLMs may not consistently follow instructions to generate a valid virtual attendee ID for the next turn to speak. Instead, it tends to predict the LLM agent to speak next, particularly when someone directly mentioned the LLM agent in the previous turn. To mitigate this issue, we randomly select the virtual attendee and their corresponding speaking role.
  • Thirdly, virtual attendees suffer from repeating the same dialog act (e.g. asking questions, pinging the LLM agent) for consecutive turns. This issue might be due to the nature of the generative model which focuses on predicting the next token. To address this issue, we introduce a cool-down mechanism for some dialog acts such as asking questions, pinging the LLM agent, and topic transition.
@@ 5.1 Future Work
Component Orchestration: MUCA integrates several components, enabling actions such as "participation encouragement" and "initiative summarization". These components have been carefully designed, tuned, and ranked to provide a harmonious experience to the chat participants. It can be beneficial to explore an easy plug-and-play method for users to design and incorporate new components into the framework without intensive tuning. Such a feature could be important, as different conversation scenarios may require chatbots to provide different set of functionalities.
Human-in-the-loop Feedback Iteration: Full user studies for feedback are costly and time-consuming. To continuously improve the chatbot post-launch, it is useful to collect implicit and explicit user behavior signals. This data should be easily transformable for automatic or semi-automatic chatbot enhancements.
Rapidly Advancing AI Technologies: The proposed MUCA framework is based on recent state-of-the-art LLMs, each with its unique style and best practices for prompting. It would be beneficial to investigate methods for updating the underlying AI models without the need of completely redoing prompting or component orchestration.
@@
Multi-Chatbot Design: The study concentrates on multi-user and single-chatbot interactions. However, scenarios involving interactions among multiple users and and multiple chatbots with different characteristics can merit further investigation. For instance, in cross-disciplinary meetings, chatbots could serve as hosts, minute-takers, or subject matter experts, offering insights to human participants as needed.
@@ A.1 MUCA Algorithm Full Implementation
The full algorithm is shown in Algorithm A.1, which allows immediate chime-in (direct chatting) for MUCA to fulfill participant’s request by detecting a keyword @mubot in the user’s last utterance.
@@
Figure 12: Data flow for Dialog Analyzer, which includes participant feature extractor and three LLM-based modules – sub-topic status update, utterance feature extractor, and accumulative summary update. The placeholders (printed in purple) in the prompts are filled by sub-topics from the Sub-topic Generator, conversation signals such as attendee names, and utterances in $U_{N_{sw},i}$. The generated outputs (sub-topic status, accumulative summary, and sub-topic being discussed, all printed in yellow) will be fed back to the sub-topics status update and accumulative summary update as inputs for the execution in the next round.
@@ A.2.1 Sub-topic Generator
To facilitate the sub-topic generation, users are required to provide user-input information $I$, which includes a topic and may optionally include an agenda, hints, and attendee roles. Hints represent any supplementary details users wish to emphasize, while attendee roles refer to the positions designated to participants. We observed that hints and attendee roles can help reduce arbitrary and random sub-topics.
Figure 11 shows the data flow for the generator with sample prompts and prompt templates. The purple text represents placeholders and is replaced by user-input Topic, Hints, and Attendee Role during execution to generate 3 sub-topics. Note that Agenda is optional and is not shown in the figure. Sub-topic Generator executes once before the chat starts, and the sub-topics remain the same throughout the conversation.
Figure 13: Data flow for direct chatting dialog act, which includes the user-input message, its full prompting, and MUCA’s responses. The placeholders (printed in purple) in the prompt are filled with user-input topic, previous accumulative summary, sub-topics status, $U_{N_{sw},i}$, sub-topics being discussed, and the last utterance. The generated responses for different types of user-input messages corresponding to different rules in the prompting are depicted.
@@ A.2.2 Dialog Analyzer
Figure 12 shows the data flow for the Dialog Analyzer. Only the participants feature extractor sub-module is based on statistical computation and the rest of the three sub-modules (sub-topic status update, utterance feature extractor, and accumulative summary update) are based on LLM inference results. Complete input prompt templates for the three LLM-based sub-modules where the purple and yellow text represent placeholders are shown. The purple ones are replaced by sub-topics from the sub-topic generator, conversation signals such as attendee names and utterances in the current context window, and the yellow ones are replaced by the generated outputs (sub-topic status, summary, and current sub-topic) from other modules. The outputs of the Dialog Analyzer will be fed into the downstream Utterance Strategies Arbitrator module to select the suitable dialog act for the response generation.
@@
Sub-topic Status Update: Three statuses are defined (i.e., not discussed, being discussed, and well discussed), which are defined differently for different types of topics to avoid ambiguity. For example, for goal-oriented communication tasks (i.e., decision-making, problem-solving, and open discussion), well-discussed refers to a situation where a consensus has been reached by the majority of participants, and the remaining participants do not have any conflicting suggestions. However, in the context of chit-chat communication, well-discussed implies that participants engage in conversations closely related and relevant to the sub-topic at hand. Even though we focus on goal-oriented communication tasks (i.e., estimation, decision-making, problem-solving, and open discussion) in our study, we show an example of chit-chat conversation in Appendix. A.6.
@@
We have observed that differentiating between being discussed and well discussed tends to be challenging, despite providing definitions for each term to the LLMs. To improve the precision and accurately update the status, the prompting utilizes the CoT style (as shown in Figure 12). This method initially returns the sub-topic IDs, followed by summarizing each participant’s opinion using generated IDs based on the current context window and previous summary. Subsequently, it determines whether the discussion meets the definition of well discussed according to the summary. Finally, the status is updated based on the results.
@@
Figure 14: Data flow for sub-topic transition dialog act, which includes hints for different trigger conditions, its full prompts, and MUCA’s responses. The placeholders (printed in purple) in the prompt are filled by the current sub-topic, next sub-topic candidates, previous summary, and $U_{N_{sw},i}$ with a size of $N_{sw}$. The generated responses corresponding to different trigger conditions are depicted.
@@
Accumulative Summary Update: This sub-module updates the summary for each participant under different sub-topics of all past utterances based on the current context window with a size of $N_{sw}$, current sub-topics being discussed from utterance feature extractor and previous summary, enabling build a memory into the MUCA system. The succinct summary is updated and stored with respect to each participant under each sub-topic, which is inspired by LLM orchestration tools like LangChain Chase (2022) that extract entities from stored messages and only returns information about entities referenced in the current run. Instead of relevant message retrieval based on the user-input query to generate a response, we provide a succinct summary for downstream modules to save computation costs for query understanding and retrieval.
@@ A.2.3 Utterance Strategies Arbitrator
Direct Chatting: In Figure 13, we show that the MUCA responds adaptively to different types of requests: greeting/acknowledgment, out-of-scope request, statement/comments, and regular request. To avoid hallucination, we disallow MUCA to make invalid assumptions beyond the user-input information, but reasonable reasoning based on chatting history is allowed (See the second example in Figure 13).
@@
Figure 16: Data flow for participation encouragement dialog act, which includes customized inactive participant status (degree of interest), its full prompting, and MUCA’s responses. The placeholders (printed in purple) in the prompt are filled with customized inactive participant status, previous accumulative summary from the upstream module, and $U_{N_{sw},i}$. The generated responses corresponding to different inactive participant statuses are depicted.
Figure 17: Prompting for multi-user simulator (MUS). All placeholders (purple) are filled with pre-defined settings or outputs from upstream modules and the LLM agent when generating the input prompts for LLM. In the prompt, we use speaking personalities to replace utterance traits for better interpretation of LLM.
Participation Encouragement: Figure 15 shows that the proposed MUCA provides customized message based on the participants’ chatting statuses, which is recorded and transferred to the hints for LLM based on the frequency of each participant’s chime-in for various sub-topics.
Figure 18: Two examples of group conversations between the proposed MUS and the proposed MUCA.
Figure 19: The prompting of baseline-small. The placeholders (printed in purple) in the prompt are filled with user input information, a short context window, and attendees’ names.
@@
Sub-topics Transition: Figure 14 shows the MUCA transits sub-topic based on different trigger conditions. When condition 1 is met, MUCA will generate a hint that instructs this dialog act to ask the participants if they are still interested in the current sub-topic without proactively proposing the next sub-topic to discuss. On the other hand, when condition 2 is met, MUCA will generate a hint that instructs this dialog act to ask the participants if they are willing to consider switching to the new sub-topic that MUCA suggests.